How many ounces of pure water must be added to 75 oz of an 8% salt solution to make a 5% salt solution?
Answer by lwsshak3(6469)

How many ounces of pure water must be added to 75 oz of an 8% salt solution to make a 5% salt solution?
**
let x=ounces of pure water to be added to 75 ounces of 8% solution
x+75=ounces of 5% solution
..
8%*75=5%(x+75)
6=.05x+3.75
.05x=6-3.75=2.25
x=2.25/.05=45 ounces
ans:
45 ounces of pure water must be added to 75 oz of an 8% salt solution to make a 5% salt solution?